OpenStudy (anonymous):

DEFG is a rectangle. DF = 5x - 3and EG = x + 5. Find the value of x and the length of each diagonal. A. x = 2, DF = 7, EG = 7 B. x = 2, DF = 7, EG = 12 C. x = 1, DF = 6, EG = 6 D. x = 2, DF = 6, EG = 6

jimthompson5910 (jim_thompson5910):

In any rectangle, the diagonals are congruent (ie equal in length) So, DF = EG 5x-3 = x+5 x = ???

OpenStudy (anonymous):

...... No clue.

jimthompson5910 (jim_thompson5910):

5x-3 = x+5 5x-3-x = 5 5x-x = 5+3 I'll let you finish up

OpenStudy (anonymous):

Yeah, I don't know.

jimthompson5910 (jim_thompson5910):

5x-x = 5+3 4x = 8 x = ???

OpenStudy (anonymous):

x=2?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now use this to find the length of the diagonals